# softMathematical software

### v1

@@ 4. Iterations of the self-adhesion operator
Let $\mathfrak{f}\subseteq\mathfrak{sg}$ be a CI frame closed under copying and marginalization. Then $N\mapsto\mathfrak{f}^{\diamond}(N)$ defines a frame closed under copying and marginalization. If $\mathfrak{f}$ is closed under intersection then $\mathfrak{f}^{\diamond}$ is closed under intersection as well. If $\mathfrak{f}$ is closed under lifting then $\mathfrak{f}^{\diamond}$ is closed under lifting, too.
Proof 4.2.
It is easy to see that ${\mathcal{K}}(N)\in\mathfrak{f}^{\diamond}(N)$ because ${\mathcal{K}}(O)\in\mathfrak{f}(O)$ for any $O\supseteq N$. Hence, $N\mapsto\mathfrak{f}^{\diamond}(N)$ is a frame in sense of Definition 2.1.
Consider ${\mathcal{M}}\in\mathfrak{f}^{\diamond}(N)$ and a bijection $\varphi:N\to M$. The goal is to show that ${\mathcal{R}}:=\varphi({\mathcal{M}})\in\mathfrak{f}^{\diamond}(M)$. Take $L\subseteq M$ and choose a bijection $\psi:M\to O$ such that $L=M\cap O$, $\psi_{|L}=\mbox{\rm id}_{L}$, and $(O\setminus L)\cap N=\emptyset$. By the remark n. 1 below Definition 3.3, it is enough to show that ${\mathcal{R}}$ and $\psi({\mathcal{R}})$ have an adhesion in $\mathfrak{f}(MO)$. Put $K:=\varphi_{-1}(L)\cup(O\setminus L)$ and define a bijection $\rho:N\to K$ by $\rho(i):=i$ for $i\in\varphi_{-1}(L)$ and $\rho(i):=\psi(\varphi(i))$ for $i\in N\setminus\varphi_{-1}(L)$. Since ${\mathcal{M}}$ is self-adhesive at $\varphi_{-1}(L)$, an extended model ${\mathcal{E}}\in\mathfrak{f}(NK)$ exists with ${\mathcal{E}}^{\downarrow N}={\mathcal{M}}$, ${\mathcal{E}}^{\downarrow K}=\rho({\mathcal{M}})$ and $(N\setminus\varphi_{-1}(L))\perp\!\!\!\perp(K\setminus\varphi_{-1}(L))\,|\,\varphi_{-1}(L)\,\,[{\mathcal{E}}]$. Extend $\varphi$ by putting $\varphi(i):=i$ for $i\in K\setminus N=O\setminus L$ and have $\varphi:NK\to MO$. Then $\varphi({\mathcal{E}})\in\mathfrak{f}(MO)$ by the assumption on $\mathfrak{f}$. The fact that copying commutes with marginalization together with the identity $\varphi_{|K}\circ\rho=\psi\circ\varphi_{|N}$ allows one to observe $\varphi({\mathcal{E}})^{\downarrow M}={\mathcal{R}}$, $\varphi({\mathcal{E}})^{\downarrow O}=\psi({\mathcal{R}})$ and $(M\setminus L)\perp\!\!\!\perp(O\setminus L)\,|\,L\,\,[\,\varphi({\mathcal{E}})\,]$.
@@
Assume ${\mathcal{M}},{\mathcal{R}}\in\mathfrak{f}^{\diamond}(N)$ and $\varphi:N\to M$ with $L:=N\cap\,M$ and $\varphi_{\,|L}=\mbox{\rm id}_{L}$. To show that ${\mathcal{M}}\cap{\mathcal{R}}$ is self-adhesive at $L$ one finds the respective extensions ${\mathcal{M}}\mapsto{\mathcal{E}}_{{\mathcal{M}}}$, ${\mathcal{R}}\mapsto{\mathcal{E}}_{{\mathcal{R}}}$, both in $\mathfrak{f}(NM)$ and puts ${\mathcal{E}}:={\mathcal{E}}_{{\mathcal{M}}}\cap{\mathcal{E}}_{{\mathcal{R}}}\in\mathfrak{f}(NM)$, by the assumption on $\mathfrak{f}$. Since the intersection commutes both with marginalization and copying one gets ${\mathcal{E}}^{\downarrow N}={\mathcal{M}}\cap{\mathcal{R}}$ and ${\mathcal{E}}^{\downarrow M}=\varphi({\mathcal{M}}\cap{\mathcal{R}})$. The definition of ${\mathcal{E}}$ also implies $(N\setminus M)\perp\!\!\!\perp(M\setminus N)\,|\,L\,\,[{\mathcal{E}}]$.
@@
Assume ${\mathcal{M}}\in\mathfrak{f}^{\diamond}(N)$ and, without loss of generality, $O\supseteq N$ with $O\setminus N$ being a singleton $o$. We distinguish two cases. The first one is to show that ${\mathcal{M}}_{N\nearrow O}$ is self-adhesive at $L\subseteq N$ relative to $\mathfrak{f}$. Consider a bijection $\varphi:O\to\tilde{M}$ with $L=O\cap\tilde{M}$ and $\varphi_{\,|L}=\mbox{\rm id}_{L}$; put $M:=\tilde{M}\setminus\varphi(o)$. The assumption on ${\mathcal{M}}$ implies the existence of ${\mathcal{E}}\in\mathfrak{f}(NM)$ with ${\mathcal{E}}^{\downarrow N}={\mathcal{M}}$, ${\mathcal{E}}^{\downarrow M}=\varphi({\mathcal{M}})$ and $(N\setminus M)\perp\!\!\!\perp(M\setminus N)\,|\,L\,\,[{\mathcal{E}}]$. Then $O\tilde{M}$ is a disjoint union of $NM$ and $o\varphi(o)$ and we can put ${\mathcal{R}}:={\mathcal{E}}_{NM\nearrow O\tilde{M}}$, which belongs to $\mathfrak{f}(O\tilde{M})$, by the assumption on $\mathfrak{f}$. The definition of lifting and the fact that it commutes with copying allow one to observe ${\mathcal{R}}^{\downarrow O}={\mathcal{M}}_{N\nearrow O}$, ${\mathcal{R}}^{\downarrow\tilde{M}}=\varphi({\mathcal{M}}_{N\nearrow O})$ and $(O\setminus\tilde{M})\perp\!\!\!\perp(\tilde{M}\setminus O)\,|\,L\,\,[{\mathcal{R}}]$.
@@
The second case is to show that ${\mathcal{M}}_{N\nearrow O}$ is self-adhesive at $Lo$ (for $L\subseteq N$) relative to $\mathfrak{f}$. Consider a bijection $\varphi:O\to\tilde{M}$ with $Lo:=O\cap\tilde{M}$ and $\varphi$ identical on $Lo$ ; put $M:=\tilde{M}\setminus o$. Choose an adhesion ${\mathcal{E}}\in\mathfrak{f}(NM)$ of ${\mathcal{M}}$ and $\varphi({\mathcal{M}})$ (at $L$). As $O\tilde{M}=NMo$ put ${\mathcal{R}}:={\mathcal{E}}_{NM\nearrow O\tilde{M}}\in\mathfrak{f}(O\tilde{M})$. Analogous arguments allow one to show that ${\mathcal{R}}$ is an adhesion of ${\mathcal{M}}_{N\nearrow O}$ and $\varphi({\mathcal{M}}_{N\nearrow O})$.
@@
{open}
Find out whether $\mathfrak{sg}$ or $\mathfrak{st}$ is an example of a CI frame $\mathfrak{f}\subseteq\mathfrak{sg}$ closed under tight replication such that $\mathfrak{f}^{\diamond}$ is not closed under tight replication.
Thus, we know from Lemma 4.1 that, if $\mathfrak{f}$ is a CI frame allowing to introduce self-adhesive models, then $\mathfrak{f}^{\diamond}$ is another such a frame.
@@
The self-adhesivity operator $\mathfrak{f}\mapsto\mathfrak{f}^{\diamond}$, applied to CI frames $\mathfrak{f}\subseteq\mathfrak{sg}$ closed under copying and marginalization, is
@@
Proof 4.4.
This is easy to observe on basis of Definition 3.3.
Corollary 6.
Let $\mathfrak{f}$ be a CI frame closed under copying and marginalization such that $\mathfrak{pr}\subseteq\mathfrak{f}\subseteq\mathfrak{sg}$. Then $\mathfrak{f}^{\diamond}$ has the same properties and $\mathfrak{pr}\subseteq\mathfrak{f}^{\diamond}\subseteq\mathfrak{f}$.
Proof 4.5.
Use the first claim in Lemma 4.1. Then Lemmas 2.3, 3.7 and 4.3 allow one to write $\mathfrak{pr}=\mathfrak{pr}^{\diamond}\subseteq\mathfrak{f}^{\diamond}\subseteq\mathfrak{f}$.
Thus, the new CI frame $\mathfrak{f}^{\diamond}$ yields a tighter approximation of $\mathfrak{pr}$ than the original frame $\mathfrak{f}$. If, moreover, $\mathfrak{f}$ is closed under intersection, then, by the second claim in Lemma 4.1, $\mathfrak{f}^{\diamond}$ is also closed under intersection. Therefore, by facts from Section 2.3, for any $N$, the family $\mathfrak{f}^{\diamond}(N)$ can be characterized in terms of CI implications (in the form of Horn clauses), which break down into permutational equivalence classes. By Corollary 6, each of these $\mathfrak{f}^{\diamond}$-rules
  • is a valid probabilistic CI implication, while
  • any of the $\mathfrak{f}$-rules (and, thus, any semi-graphoidal rule) is a valid $\mathfrak{f}^{\diamond}$-rule.
Additionally, if $\mathfrak{f}$ is closed under intersection then, by utilizing the algorithm in Lemma 3.4 as a subroutine, we can compute the $\mathfrak{f}^{\diamond}$-closure of any model (see Definition 2.25).
@@
Let $\mathfrak{f}$ be a CI frame closed under marginalization, copying, and intersection such that $\mathfrak{f}\subseteq\mathfrak{sg}$ and $\mathfrak{f}^{\diamond}(N)\subseteq\mathfrak{f}^{\diamond}(M)$ if $N\subseteq M$. Then the $\mathfrak{f}^{\diamond}$-closure of any given ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ can be computed by the following procedure:
@@
An ascending chain of CI models ${\mathcal{M}}_{0}\subseteq\cdots\subseteq{\mathcal{M}}_{i}\subseteq\cdots\subseteq{\mathcal{K}}(N)$ is obtained by the iteration, which must eventually stabilize, say at ${\mathcal{M}}_{k}$. Since $\mathfrak{f}^{\diamond}({\mathcal{M}})\in\mathfrak{f}(N)$ is self-adhesive at any $L\subseteq N$ relative to $\mathfrak{f}$, ${\mathcal{M}}_{i}\subseteq\mathfrak{f}^{\diamond}({\mathcal{M}})$ implies, by Definition 3.3, $\mathfrak{f}^{\diamond}({\mathcal{M}}_{i}|L)\subseteq\mathfrak{f}^{\diamond}({\mathcal{M}})$. Thus, one can show by induction that ${\mathcal{M}}_{i}\subseteq\mathfrak{f}^{\diamond}({\mathcal{M}})$ at any stage of the algorithm. Since ${\mathcal{M}}_{k}$ is a fixed point of the map ${\mathcal{M}}\mapsto\bigcup_{L\subseteq N}\mathfrak{f}^{\diamond}({\mathcal{M}}|L)$, it is self-adhesive relative to $\mathfrak{f}$ at any subset $L\subseteq N$ (simultaneously), and, thus, it equals to the $\mathfrak{f}^{\diamond}$-closure of ${\mathcal{M}}$.
@@
We wish to know whether or not the self-adhesivity operator $\mathfrak{f}\mapsto\mathfrak{f}^{\diamond}$ is idempotent, which means $\mathfrak{f}^{\diamond\diamond}=\mathfrak{f}^{\diamond}$. [^8] By Lemma 4.3, $\mathfrak{f}\mapsto\mathfrak{f}^{\diamond}$ would then be interpretable as the “interior” operation, which is a kind of counterpart of the “closure” operation discussed in Section 2.3. Since $\mathfrak{sg}$ is closed under intersection, we were able to use a refined version of the procedure from Lemma 4.6 to compute the iterated self-adhesive models $\mathfrak{sg}^{\diamond\diamond}(N)$ in case $|N|=4$. The basic idea of the algorithm is described in Appendix B. The result of our long computations was the conclusion that $\mathfrak{sg}^{\diamond\diamond}(N)=\mathfrak{sg}^{\diamond}(N)$ if $|N|=4$. An analogous question concerning the $\mathfrak{st}$-frame remains open.
{open}

### v3

@@ 5. Self-adhesion operator and its iterating
Let $\mathfrak{f}$ be a CI frame closed under copying and marginalization.
Then $N\mapsto\mathfrak{f}^{\diamond}(N)$ defines a CI frame $\mathfrak{f}^{\diamond}$ which is also closed under copying and marginalization.
  • (i)
    If $\mathfrak{f}$ is, additionally, closed under intersection, then the same holds for $\mathfrak{f}^{\diamond}$.
  • (ii)
    If $\mathfrak{f}$ is, additionally, closed under lifting, then the same holds for $\mathfrak{f}^{\diamond}$.
  • (iii)
    If $\mathfrak{f}$ is, additionally, closed under intersection, lifting, and ascetic extension, then so is $\mathfrak{f}^{\diamond}$.
Proof.
It is easy to see that ${\mathcal{K}}(N)\in\mathfrak{f}^{\diamond}(N)$ because ${\mathcal{K}}(O)\in\mathfrak{f}(O)$ for any $O\supseteq N$. Hence, $N\mapsto\mathfrak{f}^{\diamond}(N)$ is an abstract CI frame in sense of Definition 2.1.
Consider ${\mathcal{M}}\in\mathfrak{f}^{\diamond}(N)$ and a bijection $\varphi:N\to M$. The goal is to show that ${\mathcal{R}}:=\varphi({\mathcal{M}})\in\mathfrak{f}^{\diamond}(M)$. Take $L\subseteq M$ and choose a bijection $\psi:M\to O$ such that $L=M\cap O$, $\psi_{|L}=\mbox{\rm id}_{L}$, and ${(O\setminus L)\cap N=\emptyset}$. By the remark n. 1 below Definition 4.3, it is enough to show that ${\mathcal{R}}$ and $\psi({\mathcal{R}})$ have an adhesion in $\mathfrak{f}(MO)$. Put $K:=\varphi_{-1}(L)\cup(O\setminus L)$ and define a bijection $\rho:N\to K$ by $\rho(i):=i$ for $i\in\varphi_{-1}(L)$ and $\rho(i):=\psi(\varphi(i))$ for $i\in N\setminus\varphi_{-1}(L)$. Since ${\mathcal{M}}$ is self-adhesive at $\varphi_{-1}(L)$, an extended model ${\mathcal{E}}\in\mathfrak{f}(NK)$ exists with ${\mathcal{E}}^{\downarrow N}={\mathcal{M}}$, ${\mathcal{E}}^{\downarrow K}=\rho({\mathcal{M}})$ and $(N\setminus\varphi_{-1}(L))\perp\!\!\!\perp(K\setminus\varphi_{-1}(L))\,|\,\varphi_{-1}(L)\,\,[{\mathcal{E}}]$. Extend $\varphi$ by putting $\varphi(i):=i$ for $i\in K\setminus N=O\setminus L=O\setminus M$ and have $\varphi:NK\to MO$. Then $\varphi({\mathcal{E}})\in\mathfrak{f}(MO)$ by the assumption on $\mathfrak{f}$. The fact that copying commutes with marginalization together with the identity $\varphi_{|K}\circ\rho=\psi\circ\varphi_{|N}$ allows one to observe $\varphi({\mathcal{E}})^{\downarrow M}={\mathcal{R}}$, $\varphi({\mathcal{E}})^{\downarrow O}=\psi({\mathcal{R}})$ and $(M\setminus L)\perp\!\!\!\perp(O\setminus L)\,|\,L\,\,[\,\varphi({\mathcal{E}})\,]$.
@@
To verify (i) assume ${\mathcal{M}},{\mathcal{R}}\in\mathfrak{f}^{\diamond}(N)$ and $\varphi:N\to M$ with $L:=N\cap\,M$ and $\varphi_{\,|L}=\mbox{\rm id}_{L}$. To show that ${\mathcal{M}}\cap{\mathcal{R}}$ is self-adhesive at $L$ one finds the respective extensions ${\mathcal{M}}\mapsto{\mathcal{E}}_{{\mathcal{M}}}$, ${\mathcal{R}}\mapsto{\mathcal{E}}_{{\mathcal{R}}}$, both in $\mathfrak{f}(NM)$ and puts ${\mathcal{E}}:={\mathcal{E}}_{{\mathcal{M}}}\cap{\mathcal{E}}_{{\mathcal{R}}}\in\mathfrak{f}(NM)$, by the assumption on $\mathfrak{f}$. Since the intersection commutes both with marginalization and copying one gets ${\mathcal{E}}^{\downarrow N}={\mathcal{M}}\cap{\mathcal{R}}$ and ${\mathcal{E}}^{\downarrow M}=\varphi({\mathcal{M}}\cap{\mathcal{R}})$. The definition of ${\mathcal{E}}$ also implies $(N\setminus M)\perp\!\!\!\perp(M\setminus N)\,|\,L\,\,[{\mathcal{E}}]$.
@@
To verify (ii) assume ${\mathcal{M}}\in\mathfrak{f}^{\diamond}(N)$ and, without loss of generality, $O\supseteq N$ with $O\setminus N$ being a singleton $o$. We distinguish two cases. The first one is to show that ${\mathcal{M}}_{N\nearrow O}$ is self-adhesive at $L\subseteq N$ relative to $\mathfrak{f}$. Consider a bijection $\varphi:O\to\widetilde{M}$ with $L=O\cap\widetilde{M}$ and $\varphi_{\,|L}=\mbox{\rm id}_{L}$; put $M:=\widetilde{M}\setminus\varphi(o)$. The assumption on ${\mathcal{M}}$ implies the existence of ${\mathcal{E}}\in\mathfrak{f}(NM)$ with ${\mathcal{E}}^{\downarrow N}={\mathcal{M}}$, ${\mathcal{E}}^{\downarrow M}=\varphi({\mathcal{M}})$ and $(N\setminus M)\perp\!\!\!\perp(M\setminus N)\,|\,L\,\,[{\mathcal{E}}]$. Then $O\widetilde{M}$ is a disjoint union of $NM$ and $o\varphi(o)$ and we can put ${\mathcal{R}}:={\mathcal{E}}_{NM\nearrow O\widetilde{M}}$, which belongs to $\mathfrak{f}(O\widetilde{M})$, by the assumption on $\mathfrak{f}$. The definition of lifting and the fact that it commutes with copying allow one to observe ${\mathcal{R}}^{\downarrow O}={\mathcal{M}}_{N\nearrow O}$, ${\mathcal{R}}^{\downarrow\widetilde{M}}=\varphi({\mathcal{M}}_{N\nearrow O})$ and $(O\setminus\widetilde{M})\perp\!\!\!\perp(\widetilde{M}\setminus O)\,|\,L\,\,[{\mathcal{R}}]$.
@@
The second case is to show that ${\mathcal{M}}_{N\nearrow O}$ is self-adhesive at $Lo$ (for $L\subseteq N$) relative to $\mathfrak{f}$. Consider a bijection $\varphi:O\to\widetilde{M}$ with $Lo:=O\cap\widetilde{M}$ and $\varphi$ identical on $Lo$ ; put $M:=\widetilde{M}\setminus o$. Choose an adhesion ${\mathcal{E}}\in\mathfrak{f}(NM)$ of ${\mathcal{M}}$ and $\varphi({\mathcal{M}})$ (at $L$). As $O\widetilde{M}=NMo$ put ${\mathcal{R}}:={\mathcal{E}}_{NM\nearrow O\widetilde{M}}\in\mathfrak{f}(O\widetilde{M})$. Analogous arguments allow one to show that ${\mathcal{R}}$ is an adhesion of ${\mathcal{M}}_{N\nearrow O}$ and $\varphi({\mathcal{M}}_{N\nearrow O})$.
The verification of (iii) means, owing to (i) and (ii), to show that, given $N\subseteq M$, one has $\mathfrak{f}^{\diamond}(N)\subseteq\mathfrak{f}^{\diamond}(M)$. Since, however, we already know that $\mathfrak{f}^{\diamond}$ is closed under extension and intersection, it reduces to showing that ${\mathcal{K}}(N)\in\mathfrak{f}^{\diamond}(M)$ (see the comments below Definition 2.4). Thus, we put ${\mathcal{M}}:={\mathcal{K}}(N)$ and regard ${\mathcal{M}}$ as a CI model over $M$. We know that ${\mathcal{M}}\in\mathfrak{f}(M)$ by the assumption on $\mathfrak{f}$. The goal is to verify that ${\mathcal{M}}\in\mathfrak{f}^{\diamond}(M)$. To this end consider $L\subseteq M$ and a bijection $\varphi:M\to O$ with $L=M\cap O$ and $\varphi_{\,|L}=\mbox{\rm id}_{L}$. Let us put ${\mathcal{E}}_{1}:={\mathcal{M}}_{M\nearrow MO}$ and note that ${\mathcal{E}}_{1}\in\mathfrak{f}(MO)$ by the assumption of that $\mathfrak{f}$ is closed under lifting. We also have ${\mathcal{E}}_{1}^{\downarrow M}={\mathcal{M}}$ and $(M\setminus O)\perp\!\!\!\perp(O\setminus M)\,|\,L\,\,[{\mathcal{E}}_{1}]$. A special property of ${\mathcal{M}}$, namely $ij|K\in{\mathcal{M}},~{}\widetilde{K}\subseteq K~{}\Rightarrow~{}ij|\widetilde{K}\in{\mathcal{M}}$, together with the definition of lifting, allows one to derive that $\varphi({\mathcal{M}})\subseteq{\mathcal{E}}_{1}$. Analogously, we put ${\mathcal{E}}_{2}:=\varphi({\mathcal{M}})_{O\nearrow MO}$ and observe ${\mathcal{E}}_{2}^{\downarrow O}=\varphi({\mathcal{M}})$, $(M\setminus O)\perp\!\!\!\perp(O\setminus M)\,|\,L\,\,[{\mathcal{E}}_{2}]$ and ${\mathcal{M}}\subseteq{\mathcal{E}}_{2}$. Since $\mathfrak{f}$ is closed under intersection, ${\mathcal{E}}:={\mathcal{E}}_{1}\cap{\mathcal{E}}_{2}\in\mathfrak{f}(MO)$ is an adhesion of ${\mathcal{M}}$ and $\varphi({\mathcal{M}})$. ∎
@@
Open problem 1.
We would like to know whether either $\mathfrak{sg}$ or $\mathfrak{st}$ is an example of a CI frame $\mathfrak{f}$ closed under tight replication whose self-adhesive shrinkage $\mathfrak{f}^{\diamond}$ is not closed under tight replication.
Thus, we know from the basic claim in Lemma 5.1 that, if $\mathfrak{f}$ is a CI frame allowing to introduce self-adhesive models, then $\mathfrak{f}^{\diamond}$ is another such a frame.
@@
The self-adhesivity operator $\mathfrak{f}\mapsto\mathfrak{f}^{\diamond}$, applied to CI frames $\mathfrak{f}$ closed under copying and marginalization, is
@@
Proof.
This is easy to observe on basis of Definition 4.3. ∎
Corollary 5.3.
Let $\mathfrak{f}$ be a CI frame closed under copying and marginalization such that $\mathfrak{pr}\subseteq\mathfrak{f}$. Then $\mathfrak{f}^{\diamond}$ has the same properties and $\mathfrak{pr}\subseteq\mathfrak{f}^{\diamond}\subseteq\mathfrak{f}$.
Proof.
Use the basic claim in Lemma 5.1. Then write $\mathfrak{pr}=\mathfrak{pr}^{\diamond}\subseteq\mathfrak{f}^{\diamond}\subseteq\mathfrak{f}$ by Lemmas 2.5, 4.7 and 5.2. ∎
Thus, the new CI frame $\mathfrak{f}^{\diamond}$ yields a tighter approximation of $\mathfrak{pr}$ than the original frame $\mathfrak{f}$. If, moreover, $\mathfrak{f}$ is closed under intersection, then, by Lemma 5.1(i), $\mathfrak{f}^{\diamond}$ is also closed under intersection. Therefore, by observations from Section 2.2, for any $N$, the family $\mathfrak{f}^{\diamond}(N)$ can be characterized in terms of CI implications, namely the elements of the respective canonical implicational basis (see Definition 2.14). Elements of the canonical basis break down into permutational equivalence classes and each of them is equivalent to a collections of so-called Horn clauses [21], which have singletons as consequents. By Corollary 5.3, each of these valid implications for $\mathfrak{f}^{\diamond}(N)$ (see Section 2.3)
  • is a valid probabilistic CI implication, that is, valid implication for $\mathfrak{pr}(N)$, while
  • any valid implication for $\mathfrak{f}(N)$ remains a valid implication for $\mathfrak{f}^{\diamond}(N)$.
If additional assumptions on $\mathfrak{f}$ are accepted then, by utilizing the algorithm in Lemma 4.5 as a subroutine, we can compute the $\mathfrak{f}^{\diamond}$-closure of any model (see Definition 2.20).
@@
Let $\mathfrak{f}$ be a CI frame closed under copying, marginalization, intersection, lifting, and ascetic extension. Then the $\mathfrak{f}^{\diamond}$-closure of any given ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ can be computed by the following procedure:
@@
Note that, by Lemma 5.1(iii), $\mathfrak{f}^{\diamond}$ is also closed under intersection and ascetic extension and the concept of $\mathfrak{f}^{\diamond}$-closure makes sense (see Definition 2.20). An ascending chain of CI models ${\mathcal{M}}_{0}\subseteq\cdots\subseteq{\mathcal{M}}_{i}\subseteq\cdots\subseteq{\mathcal{K}}(N)$ is obtained by the iteration, which must eventually stabilize, say at ${\mathcal{M}}_{k}$. Since $\mathfrak{f}^{\diamond}({\mathcal{M}})\in\mathfrak{f}(N)$ is self-adhesive at any $L\subseteq N$ relative to $\mathfrak{f}$, ${\mathcal{M}}_{i}\subseteq\mathfrak{f}^{\diamond}({\mathcal{M}})$ implies, by Definition 4.3, $\mathfrak{f}^{\diamond}({\mathcal{M}}_{i}|L)\subseteq\mathfrak{f}^{\diamond}({\mathcal{M}})$. Thus, one can show by induction that ${\mathcal{M}}_{i}\subseteq\mathfrak{f}^{\diamond}({\mathcal{M}})$ at any stage of the algorithm. Since ${\mathcal{M}}_{k}$ is a fixed point of the map ${\mathcal{M}}\mapsto\bigcup_{L\subseteq N}\mathfrak{f}^{\diamond}({\mathcal{M}}|L)$, it is self-adhesive relative to $\mathfrak{f}$ at any subset $L\subseteq N$ (simultaneously), and, thus, it equals to the $\mathfrak{f}^{\diamond}$-closure of ${\mathcal{M}}$. ∎
The computation of ${\mathcal{M}}_{i+1}$ in step 2 proceeds via Lemma 4.5 and requires one $\mathfrak{f}$-closure computation over a ground set of size $2|N|-|L|$ for each $L\subseteq N$. We cannot provide an upper bound on the number of iterations until ${\mathcal{M}}_{i}$ stabilizes. Matúš [27] gave an example of a model for which the computation of its $\mathfrak{sg}(N)$-closure using a similar iterative procedure requires an exponential number of iterations in $|N|$. There is no reason to expect the algorithm in Lemma 5.4 to have better worst-case complexity.
@@
We wish to know whether or not the self-adhesivity operator $\mathfrak{f}\mapsto\mathfrak{f}^{\diamond}$ is idempotent, which means $\mathfrak{f}^{\diamond\diamond}=\mathfrak{f}^{\diamond}$. [^6] By Lemma 5.2, $\mathfrak{f}\mapsto\mathfrak{f}^{\diamond}$ would then be interpretable as the “interior” operation, which is a kind of counterpart of the “closure” operation discussed in Section 2.2. Since $\mathfrak{sg}$ is closed under intersection, we were able to use a refined version of the procedure from Lemma 5.4 to compute the iterated self-adhesive models in $\mathfrak{sg}^{\diamond\diamond}(N)$ in case $|N|=4$. The basic idea of the algorithm is described in Appendix B. The result of our long computations was the conclusion that $\mathfrak{sg}^{\diamond\diamond}(N)=\mathfrak{sg}^{\diamond}(N)$ if $|N|=4$. An analogous question concerning the $\mathfrak{st}$-frame remains open.
Open problem 2.